# Quantum hydrodynamics of a polariton fluid: pure energy relaxation terms

### v1

@@ Abstract
Cavity polaritons, hybrid half-light half-matter excitations in quantum microcavities in the strong-coupling regime demonstrate clear signatures of quantum collective behavior, such as analogues of BEC and superfluidity at surprisingly high temperatures. The analysis of the formation of these states demands an account of the relaxation processes in the system. Although there are well-established approaches for the description of some of them, such as finite lifetime polariton, an external optical pump, and coupling with an incoherent excitonic reservoir, the treatment of pure energy relaxation in a polariton fluid still remains a puzzle. Here, based on the quantum hydrodynamics approach, we derive the corresponding equations where the energy relaxation term appears naturally. We analyze in detail how it affects the dynamics of polariton droplets and the dispersion of elementary excitations of a uniform polariton condensate. Although we focus on the case of cavity polaritons, our approach can be applied to other cases of bosonic condensates, where the processes of energy relaxation play an important role.
@@
The first term on the right-hand side corresponds to a kinetic energy of the condensed particles, the second term describes an interaction with an external potential corresponding to an optical or magnetic trap, and the last non-linear term corresponds to interatomic interactions treated within the framework of the mean-field and s-wave scattering approximations. Gross-Pitaevskii equation gives a perfect description of the evolution of a conservative system with both the number of particles $N=\int|\psi(\mathbf{r},t)|^{2}d^{2}\mathbf{r}$ and the total energy:
@@
The characteristic temperatures of the BEC phase transition for cold atomic gases lie in the nanokelvin diapason, which is related to the large mass of condensing atoms, such as isotopes of sodium and rubidium. This motivated the search of analogues of BECs in condensed matter systems, where effective masses of various bosonic quasiparticles are several orders of magnitude smaller, and one can expect formation of quantum collective states at much higher temperatures.
One of the attractive possibilities is the use of ensembles of exciton polaritons (also known as cavity polaritons), which appear when a strong coupling regime between excitonic and photonic modes is realized in an optical microresonator [2, 3]. Being composite half-light-half-matter particles, polaritons inherit an extremely small effective mass (about $10^{-5}$ of the mass of free electrons) and large coherence length (in the mm scale) [4] from their photonic component. However, the presence of an excitonic component results in efficient polariton-polariton interactions, which lead to an extremely strong nonlinear optical response. The polariton BEC and superfluidity were experimentally observed at surprisingly high temperatures under both optical [5, 6, 7, 8] and electrical [9] excitation.
@@
These differences require substantial modifications of the dynamic equation 1. The finite lifetime of polaritons and an external coherent pump can also be introduced straightforwardly as a simple linear decay term and a complex time exponent source on the right-hand side of the equation 1 [10]. The incoherent pump and the coupling between polaritons and incoherent excitons are usually described in the framework of the Wouters-Carusotto model, which was first formulated for scaler polaritons [11] and then generalized for the spinor case [12].
Regarding the relaxation of energy within the condensate, there have been attempts, following the original approach of Pitaevskii [13], to include it in the phenomenological model by making the Hamiltonian non-Hermitian as $\hat{H}\rightarrow(1-\Lambda)\hat{H}$, where $\Lambda$ is a small dimensionless parameter that controls the strength of dissipation [14]. However, it can be easily seen that this energy-dependent damping does not conserve the total number of particles in the condensate and therefore cannot be viewed as pure energy relaxation.
The present paper represents an attempt to construct the theory of pure energy relaxation in a polariton system. We base our consideration on the quantum hydrodynamics approach, describing the system dynamics via set of classical field Hamilton equations for canonically conjugated variables of concentration and phase two and introducing relaxation in a natural way by adding the gradient term. We analyze how this term affects the dynamics of the polariton droplets and dispersion of the elementary excitations. We consider the scalar case, focusing on pure energy relaxation only, leaving the spinor case and a description of polarization and spin relaxation for follow-up work.
@@
[…]
@@
On the other hand, energy relaxation can be introduced in Eq. 9 via additional term proportional to $-\delta\mathscr{H}/\delta\theta$, which drags the system towards the state of the minimal energy, so that the modified dynamic equations read:
@@
Dispersion of elementary excitations and superfluidity. It is a known fact that the dispersion of elementary excitations of a conservative spatially homogeneous condensate with density $\rho_{0}$ is gapless and linear in $\mathbf{k}$, and at small momenta is described by the widely known Bogoliubov formula $\omega(k)=\sqrt{g\rho_{0}/m}k$, which according to the Landau criterium corresponds to the onset of superfluidity in the system, with
@@
All stationary solutions of GPE are also solutions of (14). This follows from the fact that for that solution $\nabla\bf{j}=0$ and the term accounting for the relaxation vanishes. However, relaxation does affect the stability of solutions. To analyze this, one should use Eq. 14 the ansatz corresponding to a spatially uniform condensate propagates along the $x$ axis with the wavevector of the absolute value $\kappa$,
@@
Note that for $\kappa=0$ Eqs.17 and 19 are equivalent. However, differently from Eq.17 the imaginary part of Eq.19 can become positive, resulting in the development of the flow instability. The typical dependence of the instability increment (positive imaginary part of (17)) for an unstable condensate is shown in Fig. 3(a) as a function of $q_{x}$, $q_{y}$.
@@
The dependency of the maximum instability increment as a function of $\rho_{0}$ and $\kappa$ is shown in Fig. 3(b). As expected, it is higher for fast-moving condensates. However, the dependency on the condensate density $\rho_{0}$ at fixed $\kappa$ is not monotonous: it is always zero for $\rho_{0}=0$, then grows, reaches its maximum, and then at some density is rolled to zero so that the condensate stabilizes. It happens exactly at the point where the velocity of the condensate becomes equal to the critical velocity defined by the Landau criterion. More details on the development of the instability are given in Supp. Mat.
@@
[…]
@@
As expected, the deceleration rate increases with the condensate densities; see Supp. Mat. We have checked that at short propagation times the velocity decreased exponentially, $v=v_{0}\mathrm{exp}(-\gamma t)$ with the decay rate $\gamma$ being independent on $k_{0}$ and scaling linearly with $\rho_{m}$. At longer propagation distances the pulse broadens, its intensity decreases, and the velocity scaling with time becomes polynomial, $v\sim t^{-2}$. Polartion-polariton interactions are repelling, so that they speed up the condensate spreading and thus suppress its deceleration. A detailed discussion of these effects will be presented elsewhere.

### v2

@@ Abstract
Cavity polaritons, hybrid half-light half-matter excitations in quantum microcavities in the strong-coupling regime demonstrate clear signatures of quantum collective behavior, such as analogues of Bose-Einstein condensate and superfluidity at remarkably high temperatures. The analysis of the formation of these states demands an account of the relaxation processes in the system. Although there are well-established approaches for the description of some of them, such as finite lifetime polariton, an external optical pump, and coupling with an incoherent excitonic reservoir, the treatment of pure energy relaxation in a polariton fluid still remains a puzzle. Here, based on the quantum hydrodynamics approach, we derive the corresponding equations where the energy relaxation term appears naturally. We analyze in detail how it affects the dynamics of polariton droplets and the dispersion of elementary excitations of a uniform polariton condensate. Although we focus on the case of cavity polaritons, our approach can be applied to other cases of bosonic condensates, where the processes of energy relaxation play an important role.
@@
Here, $\psi(\mathbf{r},t)$ is the macroscopic wavefunction of the condensate, $\hbar$ is the reduced Planck constant, $m$ is the mass of the particles, $U(\mathbf{r})$ is the external trapping potential, and $g$ is the interaction coefficient proportional to the $s$-wave scattering length. The first term on the right-hand side corresponds to a kinetic energy of the condensed particles, the second term describes an interaction with an external potential corresponding to an optical or magnetic trap, and the last non-linear term corresponds to interatomic interactions treated within the framework of the mean-field and s-wave scattering approximations. Gross-Pitaevskii equation gives a perfect description of the evolution of a conservative system with both the number of particles $N=\int|\psi(\mathbf{r},t)|^{2}d^{2}\mathbf{r}$ and the total energy:
@@
The characteristic temperatures of the BEC phase transition for cold atomic gases lie in the nanokelvin temperature range, which is related to the large mass of condensing atoms, such as isotopes of sodium and rubidium. This motivated the search of analogues of BECs in condensed matter systems, where effective masses of various bosonic quasiparticles are several orders of magnitude smaller, and one can expect formation of quantum collective states at much higher temperatures.
One of the attractive possibilities is the use of ensembles of exciton polaritons (also known as cavity polaritons), which appear when a strong coupling regime between excitonic and photonic modes is realized in an optical microresonator [2, 3]. Being composite half-light half-matter particles, polaritons inherit an extremely small effective mass (about $10^{-5}$ of the mass of free electrons) and large coherence length (in the mm scale) [4] from their photonic component. However, the presence of an excitonic component results in efficient polariton-polariton interactions, which lead to a strong nonlinear optical response. The polariton Bose-Einstein condensate and superfluidity have been experimentally observed at remarkably high temperatures under optical excitation [5, 6, 7, 8]. Polariton lasing, i.e., the generation of coherent polaritons in semiconductor cavities excited by electric current, has also been reported [9].
@@
These differences require substantial modifications of the dynamic equation 1. The finite lifetime of polaritons and an external coherent pump can also be introduced straightforwardly as a simple linear decay term and a complex time exponent source on the right-hand side of the equation 1 [10]. The incoherent pump and the coupling between polaritons and incoherent excitons are usually described in the framework of the Wouters-Carusotto model, which was first formulated for scalar polaritons [11] and then generalized for the spinor case [12].
The presence of dissipation requires an external pump that creates new particles, and the steady state is determined by the balance between gain and loss. From a dynamical perspective, this implies the existence of attractors in the system. The properties of dissipative condensates can differ significantly from those of their conservative counterparts. For example, the dispersion of bogolons in a dissipative condensate is complex, with eigenenergies acquiring nonzero imaginary parts. This highlights why the dynamics of dissipative condensates represent a large and rapidly developing area of research.
Beyond gain and loss mechanisms, another crucial aspect of polariton condensate dynamics is energy relaxation, which plays an important role in the formation of steady states and the redistribution of momentum. It has been studied extensively, with models incorporating coupling to thermal reservoirs and energy-dependent gain [13, 14]. Another approach, following the original idea of Pitaevskii [15], incorporates dissipation phenomenologically by making the Hamiltonian non-Hermitian as $\hat{H}\rightarrow(1-\Lambda)\hat{H}$, where $\Lambda$ is a small dimensionless parameter characterizing the dissipation strength [16]. However, a closer inspection shows that such energy-dependent damping alters both the amplitude and the phase of the macroscopic wavefunction, leading to a non-conservation of the total number of particles. Therefore, this method cannot be interpreted as a pure energy relaxation mechanism, as it simultaneously induces particle loss. A consistent description of energy relaxation in a closed condensate system requires dissipation terms that reduce the energy while preserving the norm of the wavefunction.
The present paper represents an attempt to construct the theory of pure energy relaxation in a polariton system. Our analysis is based on the quantum hydrodynamics approach, describing the system dynamics via set of classical field Hamilton equations for canonically conjugated variables of concentration and phase two and introducing relaxation in a natural way by adding the gradient term. We analyze how this term affects the dynamics of the polariton droplets and dispersion of the elementary excitations. We consider the scalar case, focusing on pure energy relaxation only, leaving the spinor case and a description of polarization and spin relaxation for follow-up work.
@@
Dynamic equations. Let us note that the conservative Gross-Pitraevskii equation 1 is nothing but an equation for a classical field, which can be obtained using the least action $\delta S=0$, $S=\int\mathscr{L}d^{2}\mathbf{r}dt$ with the Lagrangian being (we take $U=0$ and the spatial dimensionality two characteristic for polariton systems):
@@
Energy relaxation, which occurs due to various physical processes—such as the interaction of polaritons with phonons—can only be incorporated through additional terms in Eq. 9. A natural way to phenomenologically introduce energy relaxation is by adding the term $\delta\mathscr{H}/\delta\theta$ to Eq. 9, which drives the system toward a state of minimal energy. The resulting modified dynamical equations take the form:
@@
Dispersion of elementary excitations and superfluidity. It is a known fact that the dispersion of elementary excitations of a conservative spatially homogeneous condensate with density $\rho_{0}$ is gapless and linear in $\mathbf{k}$, and at small momenta is described by the widely known Bogoliubov formula $\omega(k)=\sqrt{g\rho_{0}/m}k$, which according to the Landau criterion corresponds to the onset of superfluidity in the system, with
@@
All stationary solutions of GPE are also solutions of (14). This follows from the fact that for that solution $\nabla\bf{j}=0$ and the term accounting for the relaxation vanishes. However, relaxation does affect the stability of solutions. To analyze this, one should substitute into Eq. 14 the ansatz corresponding to a spatially uniform condensate propagates along the $x$ axis with the wavevector of the absolute value $\kappa$,
@@
Note that for $\kappa=0$ Eqs.17 and 19 are equivalent. However, differently from Eq.17 the imaginary part of Eq. 19 can become positive, resulting in the development of the flow instability. The typical dependence of the instability increment (positive imaginary part of 17) for an unstable condensate is shown in Fig. 3(a) as a function of $q_{x}$, $q_{y}$.
@@
The dependency of the maximum instability increment as a function of $\rho_{0}$ and $\kappa$ is shown in Fig. 3(b). As expected, it is higher for fast-moving condensates. However, the dependency on the condensate density $\rho_{0}$ at fixed $\kappa$ is not monotonous: it is always zero for $\rho_{0}=0$, then grows, reaches its maximum, and then at some density is rolled to zero so that the condensate stabilizes. It happens exactly at the point where the velocity of the condensate becomes equal to the critical velocity defined by the Landau criterion. More details on the development of the instability are given in Supplementary Materials [21].
@@
We have performed a series of numerical simulations that reveal that the non-zero relaxation energy rate $\lambda\neq 0$ results in the deceleration of the condensate pulse. The dependencies of the velocities of the condensates are defined as $v=d\langle x\rangle/dt$ with $\langle x\rangle=N^{-1}\int x|\psi|^{2}dx$, $N=\int|\psi|^{2}dx$ are shown in Fig. 4 for different initial peak densities and velocities.
@@
As expected, the deceleration rate increases with the condensate densities; see Supplementary Materials [21]. We have checked that at short propagation times the velocity decreased exponentially, $v=v_{0}\mathrm{exp}(-\gamma t)$ with the decay rate $\gamma$ being independent on $k_{0}$ and scaling linearly with $\rho_{m}$. At longer propagation distances the pulse broadens, its intensity decreases, and the velocity scaling with time becomes polynomial, $v\sim t^{-2}$. Polartion-polariton interactions are repelling, so that they speed up the condensate spreading and thus suppress its deceleration. A detailed discussion of these effects will be presented elsewhere.
@@
We emphasize that the main objective of this paper is to demonstrate the importance of energy relaxation and to show that this effect can be accounted for by phenomenologically introducing a specific term into the Gross-Pitaevskii equation. The advantage of this approach is that the proposed term does not significantly complicate the mathematical model, while still capturing an effect of major physical relevance.
In this work, we considered the Gross-Pitaevskii equation with all dissipative terms omitted, except for the term representing energy relaxation. This was done intentionally to avoid masking the impact of energy relaxation by other effects, such as particle absorption and generation.
In order to accurately describe the dynamics of a real polariton condensate, multiple dissipative processes must be taken into account — including particle losses and condensation from a reservoir of incoherent excitons. One can expect that the interplay between these mechanisms and energy relaxation will strongly influence the condensate dynamics.
In this context, it is important to note that the derivation of the energy relaxation term presented in this work can be directly extended to the generalized Gross-Pitaevskii equation that incorporates a broader range of effects typical of dissipative polariton systems. We believe that including the proposed relaxation term in such generalized models will contribute to a more accurate theoretical description of experimentally observed phenomena in polariton condensates.